OpenStudy (anonymous):

If y varies inversely as x, and y = 4 as x = 8, find y for the x-value of 2.

OpenStudy (gorv):

y =k/x here k is constant of proportionality

OpenStudy (anonymous):

would it be 16

OpenStudy (gorv):

y=4 and x=8 4=k/8 k=32 y=32/x at x=2 y=32/2=16